### v1

# On the Moduli Space of Discrete Morse Functions

Abstract.
In this work, we give a combinatorial-geometric model for moduli spaces of discrete Morse functions on CW complexes. We investigate properties of moduli spaces of discrete Morse functions and their relationship to the spaces of smooth Morse functions on manifolds, as well as spaces of discrete Morse matchings on CW complexes. Moreover, we provide a similar model for the moduli space of merge trees and realise the induced merge tree and the induced barcode as continuous maps between moduli spaces.
1. Introduction
Discrete Morse theory, originally introduced by Forman [For98], is a powerful framework for investigating the simple homotopy type and other properties of CW complexes. It has shown potential to be useful for the study of persistent invariants of filtered CW complexes as they appear in topological data analysis. Other than that, discrete Morse theory has numerous applications in both, pure and applied mathematics.
Moduli spaces are most prevalent in algebraic geometry but have also been studied in other fields of mathematics, mostly other subfields of geometry and topology. The general idea of moduli spaces is to provide a topological space $M$, preferably together with a convenient way to parametrize $M$, such that points of $M$ correspond to certain mathematical objects one wants to investigate and geometric features of $M$ reflect relevant properties of said objects under investigations.
The framework we choose for the moduli-space-approach to discrete Morse functions is hyperplane arrangements in vector spaces of discrete functions. We observe that Formans definition of discrete Morse functions Definition 3.2 gives rise to the Definition 3.9 of the Morse arrangement $\mathcal{A}(X)$, a hyperplane arrangement that subdivides the vector space of discrete functions $\mathbb{R}^{X}$ on a CW complex $X$ into regions, some of which belong to Forman-equivalence classes of discrete Morse functions, and others do not contain discrete Morse functions at all. This way, the introduced moduli space of discrete Morse functions allows a parametrization indexed over the face poset $D(X)$, which helps us to investigate properties of discrete Morse functions using the combinatorial properties of the Morse arrangement $\mathcal{A}$.
In this work, we want to investigate discrete Morse theory and its connection to certain other concepts in mathematics from the viewpoint of moduli spaces.
The concepts we want to relate the moduli spaces of discrete Morse funcitons to are
  1. (1)
    The moduli spaces of smooth Morse functions on manifolds.
  2. (2)
    The moduli spaces of discrete Morse matchings on regular CW complexes.
  3. (3)
    The moduli space of merge trees.
  4. (4)
    The moduli space of barcodes.
The moduli spaces of Morse functions on manifolds have been introduced in [Cer70] and further investigated in [HW73] as subspaces of the spaces of smooth functions $C^{\infty}(M)$ on compact smooth manifolds $M$, governed by a stratification given by a certain notion of regularity for critical points and critical values. The motivation to investigate Morse functions in this way was to give a solid framework for the investigation of the pseudo-isotopoy versus isotopy question and its implications for the h-cobordism theorem. Cerf already found a map that relates certain subspaces of the space of Morse functions to spaces of discrete Morse functions Proposition 4.2 decades before discrete Morse functions had been invented by Forman. We extend this map to connected components:
Theorem 4.10.
It follows from Theorem 4.8 that Cerf’s map from Proposition 4.2 extends to a map $\eta\colon\mathcal{M}(\mathcal{N})\rightarrow\mathcal{M}_{\mathcal{N}}$, where $\mathcal{M}_{\mathcal{N}}$ denotes the CW decomposition of $M$ induced by any Morse function in the connected component $\mathcal{N}$ of $M$. Moreover, Cerf’s proof of Proposition 4.2 shows that this instance of $\eta$ is a topological submersion and compatible with the respective stratifications, too.
The moduli space of discrete Morse matchings has been introduced in [CJ05] and further investigated in [CM17] and [LS21]. We show that the moduli space of discrete Morse matchings is also canonically associated to the Morse arrangement:
Proposition 5.3.
Let $X$ be a CW complex and let $\mathcal{A}$ be the Morse arrangement on $\mathbb{R}^{X}$. Then there is a canonical embedding of posets $\mathfrak{M}(X)\subset\mathcal{L}(\mathcal{A})$.
For the moduli space of merge trees, we follow a similar approach as for the moduli space of discrete Morse functions. We model merge trees as maps from combinatorial merge trees to the real numbers and topologize using a mixture of the euclidean distance and a kind of edit distance. Definition 5.15.
Proposition 5.16.
The euclidean edit distance $\operatorname{d}$ from Definition 5.15 is a pseudo metric on the moduli space of merge trees $Mer$ and a metric on the moduli space of strict merge trees $Mer_{<}$ and the moduli space of well-branched merge trees $Mer_{wb}$.
Moreover, we show the following:
Theorem 5.17.
Let $X$ be a regular CW complex. Then the map $M\colon\mathcal{M}(X)\rightarrow Mer$ from Definition 5.8 that maps a discrete Morse function to its induced merge tree is continuous.
For the relationship to the moduli space of barcodes, we use a similar approach [^1] One can identify the space of barcodes over a given combinatorial barcode in this work with the space $\mathbb{R}^{2n}$ in [BG23, Section 4.1]. as in [BG23] for the space of barcodes and a construction for the barcode induced by a merge tree similar to the one in [${}^{+}$24]. We obtain the following result:
Proposition 5.29.
The map $B\colon Mer_{wb}\rightarrow Bar$ from Definition 5.28 is continuous.
Acknowledgements
The author would like to thank Max Planck Institute for Mathematics for the great scientific environment in which this project was conducted. Moreover, the author would like to thank Florian Kranhold and Paul Mücksch for helpful discussions about the project. Most notably the author thanks his advisor, Viktoriya Ozornova, for her advice, many helpful discussions, and detailed feedback at multiple occasions.
2. Hyperplane arrangements and their associated spaces
We review some well established notions from combinatorial geometry and topology, which will prove to be useful for our endeavor. The standard references we refer to are [${}^{+}$] and [AM17]. While most notions are standard in the literature, we adapt and extend them to the setting of this article.
Definition 2.1 (compare [${}^{+}$] and [AM17]).
A real hyperplane arrangement $\mathcal{A}$ is a finite set of hyperplanes embedded in a finite-dimensional real vector space $V$. A real hyperplane arrangement is called central if all hyperplanes contain the origin. For a central arrangement $\mathcal{A}$, the center of $\mathcal{A}$ is the subspace $\mathcal{Z}(\mathcal{A})\coloneqq\bigcap_{H\in\mathcal{A}}H\subset V$ given by the intersection of all hyperplanes of $\mathcal{A}$. A central arrangement is called essential if $\mathcal{Z}(\mathcal{A})=\{0\}$. For any central arrangement $\mathcal{A}\subset V$ we call the arrangement induced by the projection $V\rightarrow\mathcal{Z}(\mathcal{A})^{\perp}$ to the orthogonal complement of the center the essential arrangement associated to $\mathcal{A}$.
An oriented hyperplane arrangement is a real hyperplane arrangement together with a choice of an orientation class of normal vectors for each hyperplane.
We call the connected components of the complement $V\setminus\mathcal{A}$ the (open) regions of the arrangement $\mathcal{A}$. Moreover, we call intersections of half-spaces associated to $\mathcal{A}$, with at least one half-space per hyperplane chosen, faces of the hyperplane arrangement. For oriented hyperplane arrangements, each of these faces $\sigma$ is associated with a sign vector $w\in\{+,-,0\}^{\mathcal{A}}$, where $+$ means that $\sigma$ lies on the positive side of the corresponding hyperplane, -- means that $\sigma$ lies on the negative side of the corresponding hyperplane, and 00 means that $\sigma$ lies inside corresponding hyperplane. The regions correspond to sign vectors with no entry 0. The face poset $\Sigma(\mathcal{A})$ is the set of faces of $\mathcal{A}$ ordered by inclusion.
While hyperplane arrangements can be defined in arbitrary vector spaces, it is often useful to use coordinate spaces, in particular partially ordered coordinate spaces. This allows us to use equations in the ordered coordinates in order to specify the arrangements we consider.
Definition 2.2.
A partially ordered coordinate space is a real vector space $V$ together with a choice of a basis which is indexed over some partially ordered set $P$.
For convenience, we sometimes choose a specific model for partially ordered coordinate spaces:
Definition 2.3.
The category of partially ordered coordinate spaces $poVect$ consists of
  1. Objects:
    vector spaces $\operatorname{Map}(P,\mathbb{R})$ for posets $P$ together with the standard basis induced by $P$, and
  2. Morphisms:
    maps $\operatorname{Map}(Q,\mathbb{R})\rightarrow\operatorname{Map}(P,\mathbb{R})$ induced by maps $P\rightarrow Q$.
In this work, we mostly consider partially ordered coordinated spaces where the bases are indexed over either face posets of CW complexes or posets associated to merge trees or barcodes. We can characterize these three frameworks as certain types of posets, which makes partially ordered coordinate vector spaces a unifying framework for them:
Definition 2.4.
A poset $P$ is of regular CW type if it is isomorphic to the face poset $D(X)$ of a regular CW complex $X$.
A poset $P$ is of merge tree type if it
  1. (1)
    is finite,
  2. (2)
    has a unique maximum $r$,
  3. (3)
    for each $a\in P$ the set $[a,r]$ is a chain, and
  4. (4)
    each non-minimal element $a\in P$ has at least two elements $b\neq b^{\prime}$ such that $b<a$ and $b^{\prime}<a$.
A poset $P$ is of barcode type if it is a finite disjoint sum of chains of length one. We call these chains the bars of the barcode.
Remark 2.5.
In the definition above, we deviated from the standard convention in the literature to consider trees oriented away from the root. We do so because merge trees come with a preferred orientation from leaves to the root, induced by increasing filtration levels.
In order to see that Definition 2.4 models merge trees, we remark (2) and (3) model arbitrary rooted trees: the chains in (3) correspond to shortest paths between the root and other nodes of the tree. The feature that inner nodes of merge trees correspond to mergers of connected components, i.e. each inner node has at least two children, is reflected in (4).
Since merge trees can, in this fashion, be considered as both, posets and trees in the graph-theoretic sense, we use the notions from both fields interchangeably. For example, we call minimal elements leaves, non-minimal elements inner nodes, and refer to the maximum as the root.
One classic example of hyperplane arrangements, which is of particular importance for this work, is the braid arrangement.
Example 2.6.
We denote by $\mathcal{H}_{n}$ the hyperplane arrangement in $\mathbb{R}^{n}$ given by the equations $x_{i}=x_{j}$ for all $i\neq j$. We call $\mathcal{H}_{n}$ the braid arrangement $\mathcal{H}_{n}$ (see [${}^{+}$, Example 2.3.3]).
We recall the definitions of the intersection poset and the poset of flats, which are in a sense dual to each other:
Definition 2.7 ([${}^{+}$, Definition 2.1.3],[AM17, 1.3.1]).
The intersection poset $\mathcal{L}(\mathcal{A})$ of a hyperplane arrangement $\mathcal{A}$ is the set of intersections of subfamilies of $\mathcal{A}$, ordered by reverse inclusion. The poset of flats $\Pi(\mathcal{A})$ is as a set the same as the intersection poset but ordered by inclusion. The elements of the poset of flats, or the intersection poset, respectively, are called the flats of $\mathcal{A}$.
If $\mathcal{A}$ is central, then $\mathcal{L}(\mathcal{A})$ is a geometric lattice of rank $r(\mathcal{A})$. Since all arrangements in this work are central, we also refer to $\mathcal{L}(\mathcal{A})$ as the intersection lattice of $\mathcal{A}$. It is straightforward to see that $\mathcal{L}(\mathcal{A})$ always has the structure of a face poset of a simplicial complex, whereas $\Pi(\mathcal{A})$ does not necessarily satisfy this property. In particular, $\Pi(\mathcal{A})$ is never the face poset of a simplicial complex if $\mathcal{A}$ is central and $\lvert\mathcal{A}\rvert>1$.
On the other hand, $\Sigma(X)$ always has the structure of the face poset of a regular CW complex (see [AM17, 1.1.8]).
Definition 2.8 ([AM17, 1.1.9]).
We call a hyperplane arrangement simplicial if $\Sigma(\mathcal{A})\setminus\{\mathcal{Z}(\mathcal{A})\}$ has the structure of a face poset of a pure [^2] A simplicial complex is called pure if all its facets have the same dimension $d$. simplicial complex.
In fact, $\Sigma(\mathcal{A})\setminus\{\mathcal{Z}(\mathcal{A})\}$ always gives a regular CW decomposition of the sphere in $\mathcal{Z}^{\perp}$.
3. The moduli space of discrete Morse functions
Our goal is to present a framework which allows us to analyze geometric properties of and homotopies between discrete Morse functions on (regular) CW complexes. It turns out that manipulations of discrete Morse functions, like cancellations of critical cells and reordering critical cells by swapping their critical values can be geometrically realized by homotopies between discrete Morse functions, i.e. paths in the space of discrete Morse functions.
In order to define the space of discrete Morse functions on a (regular) CW complex, we consider the space of all discrete functions, which the space of discrete Morse functions will turn out to be a subspace of. As common when working with discrete Morse theory, we adopt a slight abuse of notation by not distinguishing between a CW complex and its set of cells.
Definition 3.1.
Let $X$ be a CW complex. We call the space $\mathbb{R}^{X}\coloneqq\prod\limits_{\sigma\in X}\mathbb{R_{\sigma}}$ the space of discrete functions on $X$.
Here, we notice that the coordinates of the space of discrete functions are partially ordered rather than linearly. This partial order will help us to investigate which of the discrete functions are discrete Morse functions.
For that, we recall the definition of discrete Morse functions as given by Forman:
Definition 3.2 ([For98, Definition 2.1]).
Let $X$ be a CW complex. A function $f\colon X\rightarrow\mathbb{R}$ is a discrete Morse function if for every $p$-dimensional cell $\alpha^{(p)}\in X$ the function $f$ is constant on the interior $\operatorname{int}(\alpha^{(p)})$ and
  1. (1)
    #$\#\{\beta^{(p+1)}\supset\alpha|f(\beta)\leq f(\alpha)\}\leq 1$, and
  2. (2)
    #$\#\{\gamma^{(p-1)}\subset\alpha|f(\gamma)\geq f(\alpha)\}\leq 1$.
  3. (3)
    $f(\alpha)<f(\beta)$ whenever $\alpha\subset\beta$ is not regular.
Remark 3.3.
We include the general version for non-regular CW complexes mostly for the comparison to the smooth case in Section 4. For obvious reasons, we can drop condition (3) if $X$ is regular. It is highly inconvenient to work with discrete Morse theory on non-regular CW complexes because in that case the face poset does not contain all information of the homotopy type at hand. Moreover, we make the definition for the non-regular case a bit stricter than Forman in order to simplify the definition of the induced gradient field, i.e. Morse functions are automatically strictly monotone along face inclusions of codimension higher than one.
We recall that by [For98, Lemma 2.5] for each cell $\alpha\in X$ at most one of the inequalities (1) and (2) from Definition 3.2 may actually be an equality, whereas the other one has to be a strict inequality. Therefore, for each cell $\alpha\in X$ such that (1)/(2) is an equality there is exactly one cell $\beta\supset\alpha$/$\gamma\subset\alpha$ such that $f(\beta)\leq f(\alpha)$/$f(\alpha)\leq f(\gamma)$. Hence, any discrete Morse function $f\colon X\rightarrow\mathbb{R}$ induces a partial matching $\nabla f$ on $X$ by matching $\alpha$ with $u(\alpha)\coloneqq\beta$/$d(\alpha)\coloneqq\gamma$. The unmatched cells are called critical. To be precise, $\nabla f$ is a partial matching on the Hasse diagram $D(X)$ of the face poset of $X$:
Definition 3.4.
Let $X$ be a CW complex. The Hasse diagram of the face poset of $X$ is the directed graph $D(X)$ that has vertices $V(D(X))\coloneqq X$, that is the cells of $X$, and an oriented edge $(\tau,\sigma)$ whenever $\sigma\subset\tau$ is a face relation such that there exists no $\theta\in X$ with $\sigma\subset\theta\subset\tau$ ($\sigma$ is covered by $\tau$).
Moreover, for a discrete Morse function $f$ on $X$ we define the modified Hasse diagram $D_{f}(X)$ that arises from $D(X)$ by inverting the edges matched by $f$.
Remark 3.5.
In case of a regular CW complex $X$, the directed edges correspond to face relations of codimension one. For non-regular CW complexes, edges might correspond to face relations of higher codimension, which makes finding elementary collapses via matchings harder. Furthermore, for non-regular CW complexes, the gradient paths needed for the boundary map of the Morse–Smale complex are harder to spot inside the modified Hasse diagram, since they are then not just zig-zags between any matched edges and unmatched egdges.
In the literature, there is a more specialized definition of discrete Morse functions which was first proposed by Benedetti in [Ben16]:
Definition 3.6.
Let $X$ be an abstract regular CW complex. A function $f\colon X\rightarrow\mathbb{R}$ is a Morse–Benedetti function if it satisfies:
  • Monotonicity
    For cells $\sigma\subset\tau$ we have $f(\sigma)\leq f(\tau)$.
  • Semi-injectivity
    $\lvert f^{-1}(\{z\})\rvert\leq 2$ for all $z\in\mathbb{R}$.
  • Generacity
    If $f(\sigma)=f(\tau)$, then either $\sigma\subseteq\tau$ or $\tau\subseteq\sigma$ holds.
It is straightforward to see that Morse–Benedetti functions are always discrete Morse functions but in general not the other way around.
For some purposes, it is useful to weaken Benedettis definition:
Definition 3.7.
Let $X$ be an abstract regular CW complex. We call a function $f\colon X\rightarrow\mathbb{R}$ a weak Morse–Benedetti function if it is a discrete Morse function and satisfies semi-injectivity and generacy in the sense of Definition 3.6.
The only difference between weak Morse–Benedetti functions and Morse–Benedetti functions is that matched cells of Morse–Benedetti functions have the same value, whereas for weak Morse–Benedetti functions, the higher-dimensional cell might have a strictly smaller value than the smaller-dimensional cell of the matched pair.
Remark 3.8.
One central point of this work is the interplay between abstract cell complexes and geometric cell complexes. Although discrete Morse theory excels on abstract cell complexes, we need to include geometric CW complexes for the comparison to smooth Morse theory in Section 4. Hence, on one hand we need to treat discrete Morse functions on geometric CW complexes and on the other hand we want to treat discrete Morse functions on abstract simplicial complex. We consider both points of view on regular CW complexes as a middle ground, since homotopy types of regular CW complexes uniquely determined by their face posets i.e. their associated abstract CW complexes. Nonetheless, it is important to point out that even though the face poset determines its corresponding regular CW complex up to homotopy, geometric cellular morphisms of regular CW complexes are more general than morphisms of the corresponding abstract CW complexes.
The definition of discrete Morse functions gives rise to a subdivision of the space of discrete functions induced by an oriented hyperplane arrangement in $\mathbb{R}^{X}$:
We observe that each point $p\in\mathbb{R}^{X}$ represents a discrete function $f\colon X\rightarrow\mathbb{R}$ by $f(\sigma)=p_{\sigma}$. The conditions (1) and (2) (and (3)) of Definition 3.2 impose conditions on the components of a discrete function $p\in\mathbb{R}^{X}$ whenever there is a directed edge between the corresponding cells in the Hasse diagram. That is, for the question whether $p$ represents a discrete Morse function it is relevant whether the inequality $p_{\sigma}<p_{\tau}$ holds whenever $\sigma\subset\tau$ is represented by a directed edge in $D(X)$. This information is equivalent to the question on which side of the hyperplane $H_{\sigma}^{\tau}$ in $\mathbb{R}^{X}$ given by $x_{\sigma}=x_{\tau}$ the point $p$ lies. Hence, the conditions (1) and (2) (and (3)) of Definition 3.2 can be checked by analyzing the position of $p$ relative to the hyperplane arrangement $\mathcal{A}$ given by the equations $x_{\sigma}=x_{\tau}$ for all covering face relations $\sigma\subset\tau$ in $X$. We define the following oriented hyperplane arrangement:
Definition 3.9.
Let $E$ be the set of codimension 1 face relations in $X$. We define the hyperplane arrangement $\mathcal{A}(X)\coloneqq\{H_{\sigma}^{\tau}|\sigma\subset\tau\in E\}$ where $H_{\sigma}^{\tau}$ is given by the equation $x_{\sigma}=x_{\tau}$. We observe that the assignment $(\sigma\subset\tau)\mapsto H_{\sigma}^{\tau}$ defines a bijection $E\cong\mathcal{A}$. Then let $\mathcal{L}(X)\coloneqq\{+,-,0\}^{E}$ be the sign vectors for the oriented hyperplane arrangement $\mathcal{A}^{+}(X)$ obtained by orienting $\mathcal{A}(X)$ in the following way: For each hyperplane $H_{\sigma}^{\tau}$ we define the open half-space defined by $x_{\sigma}<x_{\tau}$ to be the positive side of $H_{\sigma}^{\tau}$, the open half-space given by $x_{\sigma}>x_{\tau}$ to be the negative side of $H_{\sigma}^{\tau}$, and on $H_{\sigma}^{\tau}$ itself we have the value 0. This way, $\mathcal{L}(X)$ is the set of all covectors of the oriented hyperplane arrangement $\mathcal{A}^{+}(X)$. We call $\mathcal{A}^{+}(X)$ the Morse arrangement on $\mathbb{R}^{X}$, respectively on $X$.
For any covector $x\in\mathcal{L}(X)$ we define the sets $x_{+}\coloneqq\{e\in E|x_{e}=+\}$, $x_{-}\coloneqq\{e\in E|x_{e}=-\}$, and $x_{0}\coloneqq\{e\in E|x_{e}=0\}$.
Remark 3.10.
The index set $E$ of $\mathcal{L}(X)$ is by construction the set of edges of $D(X)$. Hence, we will use the two different points of view interchangeably without explicitly mentioning it in the notation from now on. Moreover, we drop the $+$ in the notation since there is only one orientation we are interested in. Furthermore, we will denote $\mathcal{A}(X)$ and $\mathcal{L}(X)$ just by $\mathcal{A},\mathcal{L}$, respectively, when there are not multiple complexes to be confused with.
We interpret an element $x\in\mathcal{L}$ as an equivalence class of discrete functions on some regular CW complex $X$. For two functions $p,q\in\mathbb{R}^{X}$ we define $p\sim q$ if and only if for all hyperplanes $H$, the function $p$ lies on the positive side of $H$ if and only if $q$ lies on the positive side of $H$. It turns out that this way the regions which contain discrete Morse functions are identified with their equivalence classes given by inducing the same acyclic matching. The positive region, i.e. the region on the positive side of all hyperplanes, turns out to be exactly the set of all critical discrete Morse functions on $X$. A point $p\in\mathbb{R}^{X}$ being inside a hyperplane $H_{\sigma}^{\tau}$ or on the negative side of $H_{\sigma}^{\tau}$ indicates that if $p$ represents a discrete Morse function, then the cells $\sigma$ and $\tau$ are matched by $p$. Moreover, a function $p\in\mathbb{R}$ is a discrete Morse function if and only if all functions in the same region as $p$ are.
It is immediate that the Morse arrangement $\mathcal{A}\subset\mathcal{H}_{\lvert X\rvert}$ is a subarrangement of the braid arrangement Example 2.6. Moreover, $\mathcal{A}$ is always central but never essential, i.e. the intersection of all hyperplanes of $\mathcal{A}$ always contains $0\in\mathbb{R}^{X}$ but never consists of just the origin.
We give a characterization of discrete Morse functions in the language of hyperplane arrangements:
Definition 3.11.
Let $X$ be a CW complex and let $\mathcal{A}$ be the corresponding Morse arrangement. A discrete function $p\in\mathbb{R}^{X}$ is a discrete Morse function if for every cell $\sigma\in X$
  1. (1)
    #$\#\{\tau\in X|(\tau,\sigma)\in D(X),\ p_{\tau}\leq p_{\sigma}\}\leq 1$, and
  2. (2)
    #$\#\{\tau\in X|(\sigma,\tau)\in D(X),\ p_{\sigma}\leq p_{\tau}\}\leq 1$.
  3. (3)
    $p_{\sigma}<p_{\tau}$ whenever $(\sigma,\tau)\in D(X)$ corresponds to a non-regular cover relation.
We call any region of $\mathcal{A}$ that consists of discrete Morse functions a Morse region. We call the union of all Morse regions of $\mathcal{A}$ the space of discrete Morse functions on $X$ and denote it by $\mathcal{M}(X)$. We call the unique region $R_{cr}(X)$ that only contains critical discrete Morse functions, i.e. discrete Morse functions for which every cell is critical, the critical region of $\mathcal{A}$. Moreover, we define the essential space of discrete Morse functions $\mathcal{M}_{ess}(X)$ as the intersection $\mathcal{M}_{ess}(X)\coloneqq\overline{R_{cr}(X)}\cap\mathcal{M}(X)$. We define the space of Morse–Benedetti functions $\mathcal{MB}(X)\subset\mathcal{M}(X)$ as the subspace of those discrete Morse functions that satisfy monotonicity, semi-injectivity and generacy. We define the space of weak Morse–Benedetti funtcions $\mathcal{MB}_{w}(X)\subset\mathcal{M}(X)$ as the subspace of those discrete Morse functions that satisfy semi-injectivity and generacy.
Remark 3.12.
The Morse arrangement is in general not a reflection arrangement and not simplicial. For example, the critical region of the Morse arrangement on the 2-simplex $\Delta^{2}$ has 9 faces but it would need to have only 6 faces in order to be simplicial due to the dimension of $\mathcal{Z}(\mathcal{A}(\Delta^{2}))^{\perp}$. Hence, the regular CW complex associated to the Morse arrangement, as well as its subcomplex induced by Morse regions, is in general not simplicial.
The reflection arrangement associated to $\mathcal{A}(X)$ for a regular CW complex $X$ is the braid arrangement associated to the symmetric group $\mathfrak{S}_{\lvert X\rvert}$. For arbitrary CW complexes, the story is less well behaved because one has a priori no combinatorial control over cover relations being regular.
Example 3.13.
We consider the space of discrete Morse functions on the interval $I$:
$\alpha$$\gamma$$\beta$
[…]
Figure 1. The Morse arrangement in the space of discrete functions on $I$
In order to draw the picture we made use of the fact that summands of multiples of the vector $(1,1,\dots,1)$ have no effect on whether a point in $\mathbb{R}^{X}$ corresponds to a Morse region or not. Hence, we only drew the orthogonal complement $\operatorname{span}\langle(1,1,\dots,1)\rangle^{\perp}$. That is, we actually drew the essential arrangement associated to $\mathcal{A}$.
Proposition 3.14.
Let $X$ be a regular CW complex, then the space of discrete Morse functions $\mathcal{M}(X)$ is contractible.
Proof.
We prove the slightly stronger statement that $\mathcal{M}(X)$ is a star domain with the dimension function $p_{\sigma}=\operatorname{dim}(\sigma)$ as a star point. Let $f$ be a discrete Morse function, we show that $f_{t}\coloneqq(1-t)\cdot f+t\cdot\operatorname{dim}$ is a Morse function for all $t\in[0,1]$. In order to do that, we prove that for any face relation $\sigma\subset\tau$ of codimension one where $f_{t}(\tau)\leq f_{t}(\sigma)$ holds, the inequality $f(\tau)\leq f(\sigma)$ must also hold. For $t=1$ we have $f_{1}=\operatorname{dim}$ and the statement is true. For $t<1$ we have:
$f(\tau)=\frac{f_{t}(\tau)-t\cdot(\operatorname{dim}(\sigma)+1)}{1-t}\leq\frac{f_{t}(\tau)-t\cdot\operatorname{dim}(\sigma)}{1-t}\leq\frac{f_{t}(\sigma)-t\cdot\operatorname{dim}(\sigma)}{1-t}=f(\sigma).$
This implies that $f_{t}$ is a discrete Morse function for all $t$ if $f$ is a discrete Morse function. Hence, we connected any discrete Morse function to the dimension function by a line segment of discrete Morse functions. In particular, $\mathcal{M}(X)$ is contractible.
Proposition 3.15.
Let $X$ be a regular CW complex. Then we have $\mathcal{MB}(X)\subset\overline{R_{cr}(X)}$. Moreover, $\mathcal{MB}_{w}(X)$ is dense in $\mathcal{M}(X)$ and $\mathcal{MB}(X)$ is dense in $\mathcal{M}_{ess}(X)$.
Proof.
Since Morse–Benedetti functions [^3] See Definition 3.6 are monotone, they can only lie on the positive side of or on a hyperplane $H\in\mathcal{A}(X)$ but never on the negative side. Hence, all Morse–Benedetti functions lie in the closure of the critical region $\overline{R_{cr}(X)}$.
Due to generacy, weak Morse–Benedetti functions cannot have the same value on different cells unless such cells are faces of one another. By semi-injectivity, this can only happen if the two cells are faces of codimension one, i.e. if they are matched. Hence, discrete Morse functions can only fail to be weak Morse–Benedetti functions on intersections of at least one hyperplane of $\mathcal{H}\setminus\mathcal{A}$. Thus, $\mathcal{M}(X)\setminus\mathcal{MB}_{w}(X)$ is contained in the first stratum of the stratification induced by $\mathcal{H}(X)$ which means that $\mathcal{MB}_{w}(X)$ is dense in $\mathcal{M}(X)$. It follows that $\mathcal{MB}(X)$ is dense in $\mathcal{M}_{ess}(X)$ by restricting the stratification induced by $\mathcal{H}$ to $\mathcal{M}_{ess}(X)$. ∎
Having established that the space of discrete Morse functions is homotopically trivial, we are going to investigate its combinatorial and geometrical properties. Since properties of objects are ideally investigated via morphisms, we take a look at meaningful notions of morphisms in this context.
Definition 3.16.
Let $X,Y$ be regular CW complexes. We say that a set of 0-cells $\{\sigma_{i}\}\subset X$ spans a cell $\tau\in X$ if $\{\sigma_{i}\}=(\partial\bar{\tau})_{0}$.
A continuous map $\varphi\colon X\rightarrow Y$ is called cellular if it maps the $n$-skeleton $X_{n}$ to the $n$-skeleton $Y_{n}$. A map $\varphi\colon X\rightarrow Y$ is called order-preserving if it corresponds to an order-preserving map $D(\varphi)\colon D(X)\rightarrow D(Y)$. A map $\varphi\colon X\rightarrow Y$ is called simplicial if
  • it maps 0-cells to 0-cells and
  • for any set of 0-cells $\{\sigma_{i}\}\subset X$ such that $\{\sigma_{i}\}\subset X$ spans an $n$-cell for some $n$, the set $\{\varphi(\sigma_{i})\}\subset Y$ spans a $k$-cell for some $k\leq n$.
Remark 3.17.
We remark that in Definition 3.16 cellular maps belong the the framework of of geometric CW complexes, whereas order-preserving maps and simplicial maps belong to the realm of abstract CW complexes.
There is a more general notion of equivalence of CW complexes in the literature:
Definition 3.18 ([BH04, Definition 6.32]).
Let $X$ and $X^{\prime}$ be finite CW complexes. For any cell $e\subset X$ denote by $X(e)$ the smallest subcomplex of $X$ containing $e$.
The complexes $X$ and $X^{\prime}$ are called cell equivalent if and only if there is a homotopy equivalence $h\colon X\rightarrow X^{\prime}$ with the property that there is a bijective correspondence between cells in $X$ and cells in $X^{\prime}$ such that if $e\subset X$ corresponds to $e^{\prime}\subset X^{\prime}$, then $h$ maps $X(e)$ to $X^{\prime}(e^{\prime})$ and is a homotopy equivalence of these subcomplexes.
Definition 3.19.
Let $X$ and $Y$ be CW complexes. A map $\varphi\colon X\rightarrow Y$ is called non-degenerate if it maps 0-cells to 0-cells, and induces a map $D(\varphi)\colon D(X)\rightarrow D(Y)$ that preserves cover relations, and $\varphi$ induces homotopy equivalences $X(\sigma)\simeq Y(\varphi(\sigma))$ for all $\sigma\in X$.
Remark 3.20.
It is straightforward that for regular CW complexes, cell equivalences induce simplicial equivalences. Moreover, a map $\varphi$ between regular CW complexes is non-degenerate if and only if it maps 0-cells to 0-cells and $D(\varphi)$ preserves cover relations, i.e. face relations of codimension one. In more generality, injective non-degenerate maps between CW complexes are cell equivalences onto their image.
For the context of discrete Morse theory, we prefer the point of view of simplicial maps between regular CW complexes because they focus more on the combinatorial structure at hand. Nonetheless, in the context of arbitrary CW complexes one has to work with cell equivalences because they keep track of attaching maps being irregular.
Furthermore, cell equivalences induce isomorphisms of the corresponding face posets. This way, cell equivalences and non-degenerate maps are the connecting piece between geometric and abstract morphisms of regular CW complexes.
Proposition 3.21.
A map $\varphi\colon X\rightarrow Y$ between regular CW complexes is simplicial if and only if it is cellular and order preserving.
Proof.
  1. $\Rightarrow$
    Let $\sigma\subseteq\tau$ be a face relation in $D(X)$. The vertices $\sigma_{0}$ form a subset of the vertices $\tau_{0}$. This implies $\varphi(\sigma_{0})\subseteq\varphi(\tau_{0})$. Since $\varphi$ is simplicial, it follows that $\varphi(\sigma_{0})$ spans a cell of dimension less or equal to $\operatorname{dim}\tau$. Thus, $\varphi$ is order-preserving. Moreover, the $\operatorname{dim}\sigma$-skeleton of $X$ is mapped to the $\operatorname{dim}\sigma$-skeleton of $Y$.
  2. $\Leftarrow$
    The property of $\varphi$ being cellular in particular implies that $\varphi$ preserves the 0-skeleton, i.e. maps 0-cells to 0-cells. It follows inductively that for a set of 0-cells $\{\sigma_{i}\}\subset X$, the set $\{\varphi(\sigma_{i})\}\subset Y$ spans a $k$-cell if $\{\sigma_{i}\}\subset X$ spans a $n$-cell for some $k\leq n$. Hence, $\varphi$ is simplicial.
Proposition 3.22.
Non-degenerate maps of regular CW complexes are simplicial and map $n$-cells to $n$-cells. Simplicial maps of CW complexes that map $n$-cells to $n$-cells are non-degenerate.
Proof.
Since in regular CW complexes the combinatorial structure of all cells is given by chains of inclusions of the corresponding faces, the first statement follows by induction over chains of non-trivial inclusions of faces of codimension one.
For the second statement, we observe that simplicial maps in particular map 0-simplices to 0-simplices. For the preservation of cover relations we observe that $n$-cells of regular CW complexes are characterized by chains of cover relations of their faces of length $n+1$. Since these chains need to preserve their length under simplicial maps that map $n$-cells to $n$-cells, cover relations also need to be preserved because otherwise the chains would get shorter. ∎
Lemma 3.23.
Let $\varphi\colon X\rightarrow Y$ be a non-degenerate map. Then $\varphi$ induces a map $\varphi_{*}\colon\mathbb{R}^{X}\rightarrow\mathbb{R}^{Y}$ that maps $\mathcal{A}^{+}(X)$ to $\mathcal{A}^{+}(Y)$. If $\varphi$ is injective, then $\varphi_{*}$ embeds $\mathcal{A}^{+}(X)$ into $\mathcal{A}^{+}(Y)$.
Proof.
By Proposition 3.22, $\varphi$ maps $n$-cells to $n$-cells and, therefore, induces a map $\varphi_{*}\colon\mathbb{R}^{X}\rightarrow\mathbb{R}^{Y}$. Since $\varphi$ preserves cover relations in $D(X)$ and subcomplexes up to homotopy, $\varphi$ preserves the property of cover relations being regular. Hence, $\varphi_{*}$ maps $\mathcal{A}^{+}(X)$ to $\mathcal{A}^{+}(Y)$. If $\varphi$ is injective, so is $D(\varphi)$ and it follows that $\varphi_{*}$ embeds $\mathcal{A}^{+}(X)$ into $\mathcal{A}^{+}(Y)$. ∎
Remark 3.24.
The term non-degenerate comes from the fact that the same construction with degenerate simplicial maps might lead to mapping discrete Morse functions to discrete functions with degenerate cells.
In fact, if $\varphi$ is a degenerate simplicial map, $D(\varphi)$ does not preserve cover relations. Since $D(\varphi)$ is a poset map between finite posets, the fiber $D(\varphi)^{-1}(y)$ of a point $y\in D(Y)$ is either empty or an interval in $D(X)$. Intervals in $D(X)$ correspond to subspaces in $\mathbb{R}^{X}$ which means that $\varphi_{*}$ collapses the subspace corresponding to $D(\varphi)^{-1}(y)$ to one coordinate corresponding to $y$. Geometrically, this corresponds to collapsing444in the sense of collapsing a topological space, not necessarily in the sense of simple collapses the corresponding subcomplex $U\subset X$ to a point.
In order to define an induced map on the space of discrete functions one needs to make a choice for a linear map $\tilde{\varphi}_{*}\colon\mathbb{R}^{U}\rightarrow\mathbb{R}^{y}$.
Canonical options like projections to one of the coordinates or taking the sum of the coordinates in general do not preserve the property of being a discrete Morse function. The only canonical possibilities that induce a map on the space of discrete Morse functions are the ones which ensure that the images of the collapsed subcomplexes become critical. If one wants to preserve matched pair of simplices along the collapsed subcomplex, one needs to apply more elaborate constructions that depend on the given input Morse function.
Definition 3.25.
Let $X$ be a CW complex. For $\sigma\in X$, we define
  • $\operatorname{Face}(\sigma)\coloneqq\{\tau\in X\rvert\ \tau\subset\sigma\text{arbitrary face relation}\}$,
  • $\operatorname{Coface}(\sigma)\coloneqq\{\tau\in X\rvert\ \tau\supset\sigma\text{ arbitrary face relation}\}$,
  • $\operatorname{Face}_{1}(\sigma)\coloneqq\{\tau\in X\rvert\ \tau\subset\sigma\text{ cover relation in }D(X)\}$, and
  • $\operatorname{Coface}_{1}(\sigma)\coloneqq\{\tau\in X\rvert\ \tau\supset\sigma\text{ cover relation in }D(X)\}$.
Definition 3.26.
Let $\varphi\colon X\rightarrow Y$ be either a simplicial map between regular CW complexes or a non-degenerate map between CW complexes. We define an induced map $\varphi_{*}\colon\mathbb{R}^{X}\rightarrow\mathbb{R}^{Y}$ on the spaces of discrete Morse functions inductively as follows:
  1. (1)
    for every $\tau\in Y$ such that $\lvert\varphi^{-1}(\tau)\rvert=1$, i.e. $\varphi^{-1}(\tau)=\{\sigma\}$, we define
    $\varphi(f)_{\tau}\coloneqq f_{\sigma}$
    for any $f\in\mathbb{R}^{X}.$
  2. (2)
    for any $\tau\in Y$ such that $\lvert\varphi^{-1}(\tau)\rvert\neq 1$ we proceed as follows:
    Start with the minimal dimensional such $\tau$s.
    • If $\operatorname{Face}_{1}(\tau)=\emptyset$ and $\operatorname{Coface}_{1}(\tau)=\emptyset$, then define $\varphi_{*}(f)_{\tau}\coloneqq 0.$
    • If $\operatorname{Face}_{1}(\tau)=\emptyset$ and $\operatorname{Coface}_{1}(\tau)\neq\emptyset$, then define $\varphi_{*}(f)_{\tau}\coloneqq\operatorname{min}(f_{\lvert\operatorname{Coface}_{1}(\tau)})-1$.
    • If $\operatorname{Face}_{1}(\tau)\neq\emptyset$ and $\operatorname{Coface}_{1}(\tau)=\emptyset$, then define $\varphi_{*}(f)_{\tau}\coloneqq\operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+1$.
    • If $\operatorname{Face}_{1}(\tau)\neq\emptyset$ and $\operatorname{Coface}_{1}(\tau)\neq\emptyset$, then define
      $\varphi_{*}(f)_{\tau}\coloneqq\frac{\operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+\operatorname{up}(f_{\lvert\operatorname{Coface}_{1}(\tau)})}{2}$
      for any $f\in\mathbb{R}^{X}$, where
      $\operatorname{up}(f_{\lvert\operatorname{Coface}_{1}(\tau)})\coloneqq\begin{cases}\operatorname{min}(f_{\lvert\operatorname{Coface}_{1}(\tau)})&\text{ if }f\text{ is defined on some cell of }\operatorname{Coface}_{1}(\tau).\\ \operatorname{min}(f_{\lvert\operatorname{Coface}(\tau)})&\text{ if }f\text{ is only defined on cells of }\operatorname{Coface}(\tau)\setminus\operatorname{Coface}_{1}(\tau)\\ \operatorname{max}(f_{\lvert\operatorname{Face}_{1}(\tau)})+2&\text{ if }f\text{ is not defined on any cell of }\operatorname{Coface}(\tau).\end{cases}$
      Here, we slightly abuse notation because $f_{\lvert\operatorname{Coface}_{1}(\tau)}$ might not be defined for all cells of $\operatorname{Coface}_{1}(\tau)$. If $f_{\lvert\operatorname{Coface}_{1}(\tau)}$ is only defined for some cells of $\operatorname{Coface}_{1}(\tau)$, we take the minimum on those cells.
    We apply this definition inductively over the dimension of the cells for which $\varphi_{*}$ is not yet defined.
Remark 3.27.
The set $\operatorname{Face}_{1}(\sigma)$ corresponds to the set of elements $\tau$ of $D(X)$ which are covered by $\sigma$, whereas the cells of $\operatorname{Coface}_{1}(\sigma)$ are exactly the cells $\tau$ of $X$ which cover $\sigma$. If one uses the covering relation instead of the codimension, there is a straightforward generalization to non-regular CW complexes.
Proposition 3.28.
The induced map $\varphi_{*}$ is affine-linear. Moreover, $\varphi_{*}$ is linear if and only if $\operatorname{Face}_{1}(\tau)=\emptyset$ and $\operatorname{Coface}_{1}(\tau)=\emptyset$ for all $\tau$ such that $\lvert\varphi^{-1}(\tau)\rvert\neq 1$. In particular, $\varphi_{*}$ is always linear if $\varphi$ is injective.
Proof.
The proof is straightforward from the definition. ∎
Proposition 3.29.
The map $\varphi_{*}$ preserves induced matchings if $\varphi$ is injective.
Proof.
If $\varphi$ in injective, then $X$ can be interpreted as a subcomplex of $Y$. The map $\varphi_{*}(f)$ is then just an extension of $f$ to $Y$ that is critical on $Y\setminus\varphi(X)$. By construction, we have $\varphi_{*}(f)_{\lvert\varphi(X)}=f\circ\varphi^{-1}$ and, therefore, the statement holds. ∎
Example 3.30.
We consider the following simplicial map between two regular complexes:
[…]
Figure 2. A simplicial map between regular CW complexes
We consider the following discrete Morse function $f$:
[…]
The induced discrete Morse function $\varphi_{*}(f)$ is:
[…]
Remark 3.31.
As we see in Example 3.30, the map $\varphi_{*}$ only preserves matchings where it is injective. Therefore, $\varphi_{*}(f)$ might in general induce a Morse complex with more cells than $f$ does even if $\varphi$ collapses a subcomplex of $X$.
Definition 3.32.
Let $\varphi\colon X\rightarrow Y$ be a cellular map of CW complexes. Then there is an induced map $\varphi^{*}\colon\mathbb{R}^{Y}\rightarrow\mathbb{R}^{X}$ given by $\varphi^{*}g\coloneqq g\circ\varphi$, for $g\colon Y\rightarrow\mathbb{R}$.
Proposition 3.33.
Let $X\rightarrow Y$ be an injective non-degenerate map of CW complexes. Then $\varphi^{*{-1}}(\mathcal{A}(X))\subset\mathcal{A}(Y)$ is a subarrangement. Moreover, $\varphi^{*}$ restricts to a map $\mathcal{M}(Y)\rightarrow\mathcal{M}(X)$.
Proof.
The first statement follows directly by the definition of non-degenerate maps because they preserve cover relations. The second statement follows by injectivity. ∎
4. Relationship to smooth Morse theory
In this section we want to recall the description of the space of Morse functions on a smooth manifold given by Cerf in [Cer70]. Moreover, we will explain the relationship between certain neighborhoods of Morse functions and the space of discrete Morse functions on their induced CW decompositions, inspired by [Cer70, Section 3.2].
In order to get going, we start with a description of the space of smooth Morse functions on a given manifold $M$ and its relationship to Cerf’s stratifications [Cer70] of the space of smooth functions and the space of Morse functions. Throughout this section, we assume any smooth manifold $M$ to be finite-dimensional, compact, and with a fixed Riemannian metric.
Definition 4.1 ([Cer70, Definition 1,2, and 3]).
Let $M$ be a smooth manifold. Denote by $C^{\infty}(M)$ the space of smooth functions from $M$ to $\mathbb{R}$ together with the $\mathcal{C}^{\infty}$ topology. Let $f\in C^{\infty}(M)$ be a smooth function, let $p\in M$ be a critical point of $f$, and let $a\in\mathbb{R}$ be a critical value of $f$.
The codimension of $p\in M$ is defined as $\operatorname{codim}(\operatorname{\partial_{p}(f)}\subset C^{\infty}_{0,p}(M))\coloneqq$
$\operatorname{dim}(C^{\infty}_{0,p}(M)/(\operatorname{\partial_{p}(f)}))$, where $C^{\infty}_{0,p}(M)$ denotes the ring [^5] These rings do not necessarily contain a unit. of germs of smooth functions vanishing at $p$ and $(\operatorname{\partial_{p}(f)})$ denotes the ideal generated by the partial derivatives of $f$ at the point $p$.
The codimension of $a$ is defined as $\operatorname{codim}(a)\coloneqq\lvert\{q\in f^{-1}(a)|p\text{ critical}\}\rvert-1$.
The codimension of $f$ is defined as
$\operatorname{codim}(f)\coloneqq\sum\limits_{p\in M\text{ critical}}\operatorname{codim}(p)+\sum\limits_{a\in\mathbb{R}\text{ critical}}\operatorname{codim}(a).$
Moreover, we define $\mathcal{F}\coloneqq C^{\infty}(M)$ and $\mathcal{F}^{j}\coloneqq\{f\in\mathcal{F}|\operatorname{codim}(f)=j\}$. We call $\{\mathcal{F}^{j}\}_{j\in\mathbb{N}}$ the natural stratification of $\mathcal{F}$.
We recall that for a function $f\in\mathcal{F}$ having a higher codimension than 0 means either that $f$ has degenerate critical points or multiple critical points of the same value. We denote the space of (smooth) Morse functions by $\mathcal{M}_{\text{smo}}(M)\subset\mathcal{F}$.
Next, we consider Cerf’s comparison to the space of discrete functions.
Proposition 4.2 ([Cer70, Section 3.2: Definition 5, Lemma 1]).
[…]
$\begin{split}\eta\colon\mathcal{V}&\rightarrow\mathbb{R}^{q}\\ \tilde{f}&\mapsto(\tilde{f}(\tilde{c}_{1}),\dots,\tilde{f}(\tilde{c}_{q}))\end{split}$
such that the natural stratification of $\mathcal{F}$ is the preimage of the stratification induced by the hyperplane arrangement $\mathcal{H}_{q}$ (see Example 2.6).
Our plan for this section is to identify the space $\mathbb{R}^{q}$ in Proposition 4.2 with the space of discrete functions on the CW decomposition of $M$ induced by $f$, to investigate how far the neighborhood $\mathcal{V}$ extends, and to generalize $\eta$ in order to define maps which compare the space of smooth Morse functions on $M$ to spaces of discrete Morse functions on cellular decompositions of $M$.
It is well known (see e.g. [BH04, Theorem 6.34]) that Morse functions only induce a cellular decomposition $M_{f}$ of $M$, unique up to cell equivalence, if they fulfill the Morse–Smale property. Alternatively, on can associate gradient-like Morse–Smale vector fields to non-Morse–Smale functions instead (see [Mat02, Chapter 3 and Section 4.2]). Since we want to take the point of view of the moduli space of Morse functions, we have fixed a Riemannian metric and consider the subspace $\mathcal{MS}_{\text{smo}}(M)\subset\mathcal{M}_{\text{smo}}(M)$ of Morse–Smale functions on $M$.
Definition 4.3.
Let $Z$ be either the space of Morse functions $\mathcal{M}_{\text{smo}}(M)\subset\mathcal{F}$ or the space of Morse–Smale functions $\mathcal{MS}_{smo}(M)\subset\mathcal{M}_{\text{smo}}(M)\subset\mathcal{F}$ and let $f\in Z$. We denote by $\mathcal{N}(f)\subset Z$ the connected component of $f$ in $Z$ [^6] It will be clear from the context which of the two possibilities for $Z$ will be used..
It is a classical result that the property of being a Morse function is stable, i.e. for any Morse function $f$, there is an open neighborhood $U(f)\subset C^{\infty}(M)$ such that $U(f)\subset\mathcal{M}_{\text{smo}}(M)$ (see [BH04, Corollary 5.24]). We want to investigate this phenomenon in a slightly different direction: we show that the CW complex induced by a Morse–Smale function is stable on connected components of Morse–Smale functions. As a first step towards that, we have the following theorem:
Proposition 4.4.
[…]
Before we prove this proposition, we need two lemmas.
Lemma 4.5.
Let $f_{t}\colon I\rightarrow\mathcal{M}_{smo}(M)$ be a path in the space of Morse functions [^7] Sometimes called a 1-parameter family of Morse functions.. Then
  1. (1)
    the critical points of $f_{t}$, i.e. the zeros of $Df_{t}$ depend continuously on $t$, and
  2. (2)
    for any $p\in M$ the eigenvalues of $Hf_{t}(p)$ depend continuously on $t$.
Proof.
  1. (1)
    Since $f_{t}$ depends continuously on $t$ and since for any local coordinates the entries of the Jacobian $Df_{t}(p)$ at any point $p\in M$ are the first partial derivatives of $f_{t}$ at $p$, the entries of $Df_{t}$ depend continuously on $t$ because $t$ is just a parameter for the partial derivatives. By definition, $p$ is a critical point of $f$ if and only if $\operatorname{det}Df_{t}(p)\neq 0$. The statement follows because the determinant is continuous in the entries of the matrix.
  2. (2)
    Similar to (1) the entries of the Hessian $Hf_{t}(p)$ at any point $p$ depend continuously on $t$. Since $Hf_{t}(p)$ is symmetric, all its eigenvalues are real. Since the eigenvalues of $Hf_{t}(p)$ are zeros of the characteristic polynomial $\chi_{Hf_{t}(p)}$ and the coefficients of $\chi_{Hf_{t}(p)}$ depend continuously on the entries of $Hf_{t}(p)$, the statement follows directly by [Tyr97, Theorem 3.9.1].
Lemma 4.6.
Along any path $f_{t}\colon I\rightarrow\mathcal{MS}_{smo}(M)$, the number of critical points of $f_{t}$ is constant with respect to $t$.
Proof.
Since $f_{t}$ is assumed to be Morse–Smale for any $t\in I$, each $f_{t}$ for any $t\in I$ has a finite number of isolated critical points and each $f_{t}$ induces a CW decomposition $\widetilde{M}$ with a $k$-cell for each critical point of index $k$ of $f_{t}$ of $M$, i.e. a CW complex which is homotopy equivalent to $M$. Thus, the Euler characteristic of the two agree, i.e. $\chi(M)=\chi(\widetilde{M})$. Since $\chi(\widetilde{M})$ is defined as the alternating sum over the number of $k$-cells of $\widetilde{M}$, this means that if any new critical points of $f_{t}$ appear with increasing $t$, they can only appear in pairs such that they leave the Euler characteristic unchanged.
Due to Lemma 4.5 (1) we can, without loss of generality, assume that there is a $t_{0}$ such that $f_{t_{0}}$ has two more non-degenerate critical points than $f_{t}$ for $t<t_{0}$. Let $c_{1},c_{2}$ be those two new critical points. We can chose disjoint neighborhoods $c_{1}\in U_{1}$, $c_{2}\in U_{2}$ and a small perturbation $\widetilde{f}_{t}$ such that, without loss of generality, $c_{1}$ appears at a time $t_{1}>t_{0}$ but $f_{t}$ and $\widetilde{f}_{t}$ agree outside of $U_{1}$.
Such a perturbation can be achieved as follows: Choose $\epsilon>0$ such that $B_{\epsilon}(c_{1})\subset U_{1}$ with respect to the metric on $M$ induced by the given Riemannian metric. We define the smooth function $\rho\colon[t_{0}-\delta,t_{0}+\delta]\rightarrow[t_{0}-\delta,t_{0}+\delta]$ by
$\rho(t)\coloneqq\begin{cases}\frac{t+t_{0}-\delta}{2}&\text{for }t_{0}-\delta\leq t<t_{0}\\ \frac{3t-t_{0}-\delta}{2}&\text{for }t_{0}\leq t\leq t_{0}+\delta\end{cases}.$
Then we define a perturbation
$\widetilde{f}(x,t)\coloneqq\begin{cases}f(x,t)&\text{for }x\in M\setminus B_{\epsilon}(c_{1}),\ t\in I\\ f(x,t)&\text{for }x\in B_{\epsilon}(c_{1}),\ t\in I\setminus[t_{0}-\delta,t_{0}+\delta]\\ f(x,(1-\frac{d(x,c_{1})}{\epsilon})\cdot\rho(t)+\frac{d(x,c_{1})}{\epsilon}\cdot t)&\text{for }x\in B_{\epsilon}(c_{1}),\ t\in[t_{0}-\delta,t_{0}+\delta]\end{cases},$
where the formulas are supposed to be interpreted for an appropriate choice of coordinates. Since $\widetilde{f}_{t}$ has no degenerate critical points by construction, $\widetilde{f}_{t}$ is a path of Morse functions and $\widetilde{f}_{t_{0}}$ can be arbitrarily closely approximated by a Morse–Smale function $f^{\prime}_{t_{0}}$.
But then, the function $f^{\prime}_{t_{0}}$ would induce a CW decomposition of $M$ with a different Euler characteristic which is a contradiction to the CW decomposition being homotopy equivalent to $M$. Thus, no new critical points can appear along $f_{t}$.
It can be proved analogously that along $f_{t}$ the number of critical points cannot decrease. All in all, the statement follows. ∎
Proof of Proposition 4.4.
It follows directly from Lemma 4.6 that all functions in $\mathcal{N}(f)$ have the same number of critical points. Since by Lemma 4.5 the eigenvalues of $Hf_{t}(p)$ depend continuously on $t$ for all $p\in M$, the indices of critical points cannot change along a path $f_{t}$ of Morse functions because otherwise, the intermediate value theorem implies that for some $t^{\prime}\in I$ there must have been a critical point with eigenvalue 0 which is a contradiction to $f_{t}$ being a path of Morse functions. Hence, the indices of the critical points remain unchanged along any path $f_{t}$. Thus, the statement follows. ∎
Corollary 4.7.
[…]
The result Corollary 4.7 can be extended to connected components of Morse functions if one does not fix a Riemannian metric, but allows the metric to be changed in order to make Morse functions Morse–Smale or works with choosing appropriate gradient-like flows instead. Both these approaches come with the problem that they do not give rise to a well-defined assignment of a CW complex to a Morse function because they involve non-canonical choices. This is the reason why we decided to fix a Riemannian metric at the cost of restricting to the space of Morse–Smale functions with respect to the chosen Riemannian metric. This way, we made the necessary non-canonical choice beforehand.
We continue with the investigation of the induced gradient-flows because they are responsible for the attaching maps of the CW decomposition induced by a Morse–Smale function.
To fix notation, for a Morse function $f\colon M\rightarrow\mathbb{R}$ on a Riemannian manifold $M$ with Riemannian metric $g$, we denote the gradient vector field of $f$ with respect to $g$ by $\nabla f$. Moreover, we denote by $\phi_{f}(x,\theta)$ the flow generated by $-\nabla f$. Let further $\operatorname{Cr}(f)$ denote the set of critical points of $f$.
Theorem 4.8.
[…]
Proof.
Let $f_{t}$ be a path from $f$ to $g$ in $\mathcal{MS}_{\text{smo}}(M)$. Since $t$ is a parameter with respect to differentiation, $\nabla f_{t}$ depends continuously on $t$. Since the flow $\phi_{t}(x,\theta)\colon M\times\mathbb{R}\rightarrow M$ of the gradient field $\nabla f_{t}$ is the solution to the initial value problem
$\begin{split}\frac{d}{d\theta}\phi_{t}(x,\theta)&=-(\nabla f_{t})(\phi_{t}(x,\theta))\\ \phi_{t}(x,0)&=x\end{split}$
and solutions to initial value problems depend smoothly on their initial values, the time-dependent flow $\phi_{t}(x,\theta)$ depends continuously on $t$. It is a standard trick in dynamics to replace a time-dependent flow $\phi_{t}$ by a time-independent one $\Phi$ as follows:
$\Phi\colon(M\times\mathbb{R})\times\mathbb{R}\rightarrow M\times\mathbb{R},\ \Phi((x,t),\theta)\coloneqq(\phi_{t}(x,\theta),t+\theta)$
It is a classical result that for any fixed $t$ and for any starting point $x\in M$ the limits $\lim\limits_{\theta\rightarrow\pm\infty}\phi_{t}(x,\theta)$ exist and are critical points of $f_{t}$. Since $\Phi$ is a continuous flow, this implies that the critical points which are connected by a given flow line also depend continuously on $t$. Together with Proposition 4.4, the statement follows. ∎
Corollary 4.9.
[…]
Corollary 4.10.
It follows from Theorem 4.8 that Cerf’s map from Proposition 4.2 extends to a map $\eta\colon\mathcal{M}(\mathcal{N})\rightarrow\mathcal{M}_{\mathcal{N}}$, where $\mathcal{M}_{\mathcal{N}}$ denotes the CW decomposition of $M$ induced by any Morse function in the connected component $\mathcal{N}$ of $M$. Moreover, Cerf’s proof of Proposition 4.2 shows that this instance of $\eta$ is a topological submersion and compatible with the respective stratifications, too.
Building on Corollary 4.9, we can extend the map $\eta$ from Proposition 4.2 to the connected component $\mathcal{N}(f)$ of a given Morse–Smale function $f$. By Proposition 4.2, it still follows that $\eta\colon\mathcal{N}(f)\rightarrow\mathcal{M}_{\mathcal{N}(f)}\coloneqq\mathcal{M}(M_{f})$ is a topological submersion which is compatible with the respective stratifications.
Using Cerf’s stratification of the space of Morse functions, especially the connection between different connected components given by cancellations and creations of pairs of critical points, we want to further extend $\eta$ to certain systems of connected components:
Definition 4.11.
Consider the set $\pi_{0}(\mathcal{MS}_{smo}(M))$ of connected components of the space of Morse–Smale functions on $M$. Let $\mathcal{N}_{1},\mathcal{N}_{2}\in\pi_{0}(\mathcal{MS}_{smo}(M))$ be connected components. Then we define a partial order $\leq$ on $\pi_{0}(\mathcal{MS}_{smo}(M))$ generated by $\mathcal{N}_{1}<\mathcal{N}_{2}$ if for any $f\in\mathcal{N}_{2}$ there is a $g\in\mathcal{N}_{1}$ such that $g$ arises from $f$ by a cancellation of a pair of critical points. We call $\leq$ the cancellation order on $\pi_{0}(\mathcal{MS}_{smo}(M))$.
Remark 4.12.
It is shown in [Cer70, Section 3] that if two Morse functions are related to each other by a cancellation of one pair of critical points, then all paths between them are homotopic to one which traverses the first stratum $\mathcal{F}^{1}$ in one isolated point. Hence, we can think about two connected components $\mathcal{N}_{1},\mathcal{N}_{2}$ such that $\mathcal{N}_{1}<\mathcal{N}_{2}$ as being neighbors, only separated by a codimension one stratum which corresponds to a hyperplane in $\mathcal{M}(\mathcal{N}_{2})$ via $\eta$.
Proposition 4.13.
Let $f,g\in\mathcal{MS}_{smo}(M)$ be two Morse–Smale functions such that $f$ arises from $g$ by a cancellation of one pair of critical points. Then $M_{f}$ and $M_{g}$ are simply homotopy equivalent by a simple collapse $M_{g}\rightarrow M_{f}$/ a simple extension $M_{f}\rightarrow M_{g}$ of the two cells which correspond to the canceled critical points at the corresponding sublevel complex. In particular, the map $M_{f}\rightarrow M_{g}$ is non-degenerate.
Proof.
This statement is basically a rephrased version of [Mat02, Theorem 3.34 (canceling handles: rephrased)]. In [Mat02, Theorem 3.34], building on techniques from [Mil65, Proof of Theorem 5.4] the authors argue that a cancellation of a pair of critical points induces a pair of handles in the induced handle body decomposition such that the handlebody decomposition induced by $f$ is diffeomorphic to the one induced by $g$. The statement then follows in a straightforward manner from the construction of the CW decomposition associated to a handlebody decomposition [Mat02, Theorem 4.18]. ∎
Corollary 4.14.
Let $\mathcal{N}_{1}<\mathcal{N}_{2}\in\pi_{0}(\mathcal{MS}_{smo}(M))$ be connected components which are related by cancellations of critical points. Then $\mathcal{A}(M_{\mathcal{N}_{1}})$ canonically embeds into $\mathcal{A}(M_{\mathcal{N}_{2}})$ by Lemma 3.23. Moreover, we have an induced map $\mathcal{M}(M_{f})\rightarrow\mathcal{M}(M_{g})$ given by Definition 3.26.
5. Relationship to other concepts in persistent topology
5.1. The space of discrete Morse matchings
In the literature, there is already work on a so-called space, or complex, of discrete Morse functions. In fact, this name has been used with a slight abuse of terminology since these works refer to a complex whose simplices correspond to Morse matchings rather than Morse functions.
Definition 5.1 ([CJ05, Section 2]).
Let $X$ be a simplicial complex. The complex of discrete Morse matchings $\mathfrak{M}(X)$ has as vertices cover relations in $D(X)$ and simplices sets of cover relations in $D(X)$ that correspond to acyclic matchings on $D(X)$.
It is straightforward how this notion should be generalized to arbitrary CW complexes:
Definition 5.2.
Let $X$ be a CW complex. The complex of discrete Morse matchings $\mathfrak{M}(X)$ on $X$ has vertices cover relations in $D(X)$ that correspond to regular faces and simplices sets of vertices that correspond to acyclic matchings on $D(X)$.
Since the geometric structure of the simplices, i.e. their points, do not have any obvious meaning for discrete Morse matchings, we will not distinguish between $\mathfrak{M}(X)$ and its face poset $D(\mathfrak{M}(X))$ in our notation.
Proposition 5.3.
Let $X$ be a CW complex and let $\mathcal{A}$ be the Morse arrangement on $\mathbb{R}^{X}$. Then there is a canonical embedding of posets $\mathfrak{M}(X)\subset\mathcal{L}(\mathcal{A})$.
Proof.
The minimal elements of both posets are in canonical bijection because in both cases they correspond to the cover relations in $D(X)$ that correspond to regular face relations. Then collections of cover relations that form an acyclic matching in $\mathfrak{M}(X)$ are mapped to intersections of hyperplanes in $\mathcal{L}(\mathcal{A})$ that correspond to the same acyclic matching. ∎
5.2. The space of merge trees
In order to give a model for the moduli space of merge trees, we introduce a notion of edit morphisms to relate different combinatorial merge trees to each other.
Definition 5.4.
An edit move on a poset of merge tree type $P$ is one of the following operations:
  1. The identity:
    The poset $P$ is left unchanged.
  2. Adding a leaf:
    A new minimal element $x$ is added in one of two ways: either, we just introduce the relation $x\leq y$ for a pre-existing non-minimal element $y\in P$, or we add a new non-minimal element $y^{\prime}$ between two pre-existing elements $z_{1}\leq y\leq z_{2}$ and add the relation $x\leq y$.
  3. Removing a leaf:
    A pre-existing minimal element $x\in P$ is removed. If $x$s parent node $y$ then only has one child node $x^{\prime}$, we also remove $y$.
  4. Splitting an inner node
    If a node $z\in P$ has at least 3 child nodes, $z$ can be split into two nodes $z_{1}\leq z_{2}$ such that there is no $x\in P$ with $z_{1}\leq x\leq z_{2}$. The previous child nodes of $z$ are then distributed among $z_{1}$ and $z_{2}$ such that $z_{1}$ gets at least two child nodes and $z_{2}$ gets at least one child node other than $z_{2}$.
  5. Merging two adjacent inner nodes:
    If there are two non-minimal elements $z_{1}\leq z_{2}\in P$ such that there is no element between them, i.e. there exists no $x:z_{1}\leq x\leq Z_{2}$, we construct the quotient $P/_{z_{1}=z_{2}}$. This creates an element $z$ instead of $z_{1}$ and $z_{2}$ such that all previous child nodes of $z_{1}$ and $z_{2}$ become child nodes of $z$.
An edit morphism $\eta\colon P\rightarrow P^{\prime}$ between posets of merge tree type is a composition of edit moves together with an automorphism of $P^{\prime}$. An edit morphism is called elementary if it performs only one of the edit moves. An edit morphism is called trivial if it has the identity as its only corresponding edit move.
Remark 5.5.
Let $P,P^{\prime}$ be posets of merge tree type such that $P^{\prime}$ arises from $P$ by either an addition of a leaf or a splitting of an inner node. Then any corresponding edit morphism induces a poset morphism $\eta\colon P\rightarrow P^{\prime}$ given by the inclusion of the preexisting nodes.
We identify spaces of discrete functions, as introduced in Definition 3.1, with sets of set maps $\operatorname{Map}(P,\mathbb{R})$ for posets $P$ of CW type. Moreover, we identify sets of merge trees with the same underlying combinatorial merge tree $P$ with sets of order-preserving maps $\operatorname{Map}_{\leq}(P,\mathbb{R})$ for posets $P$ of merge tree type.
Definition 5.6.
We define the set of merge trees $Mer$ as the disjoint union
$Mer\coloneqq\coprod\limits_{P\text{ of merge tree type}}\operatorname{Map}_{\leq}(P,\mathbb{R}).$
For a merge tree $\theta\in Mer$ we denote the underlying combinatorial merge tree, i.e. the source of $\theta$, by $s(\theta)$.
Moreover, we define the set of strict merge trees $Mer_{st}$ by
$Mer_{st}\coloneqq\coprod\limits_{P\text{ of merge tree type}}\operatorname{Map}_{<}(P,\mathbb{R}),$
i.e. we here demand that the height functions on the combinatorial merge trees are strictly monotone.
Furthermore, we define the set of well-branched merge trees $Mer_{wb}\subset Mer_{st}$ by
$Mer_{wb}\coloneqq\{\theta\in Mer_{st}|\forall y\in s(\theta)\text{ the restriction }\theta_{\lvert s(\theta)_{\leq y}}\text{ has a minimum.}\},$. } ,
where $s(\theta)_{\leq y}$ denotes the rooted subtree of $s(\theta)$ with root $y$.
Remark 5.7.
In order to ensure that $Mer$ is actually a set, we would need to fix a way of how to turn the category of finite sets into a small category. Once that is done, we can do the same for posets since there are only finitely many relations, in particular finitely many partial orders, on any finite set. Then posets of merge tree type form a subset of said set and the disjoint union above is actually indexed over a set.
Definition 5.8.
Let $X$ be a regular CW complex. We define a map $M\colon\mathcal{M}(X)\rightarrow Mer$ as follows: Let $f\in\mathcal{M}(X)$ be a discrete Morse function and let $\operatorname{Cr}_{i}(f)$ the set of critical cells of dimension $i$ of $f$. We create a poset of merge tree type $P$ as follows: for each critical vertex $\sigma\in\operatorname{Cr}_{0}(f)$ we add an element $\sigma$ to $P$. For each time connected components $\sigma_{1},\dots,\sigma_{k}$ of sublevel complexes merge within the filtration, we add an element $\tau$ to $P$ and add the relations $\sigma_{i}\leq\tau$. The elements $\tau$ correspond to certain critical edges $\tau\in\operatorname{Cr}_{1}(f)$. Then we define $M(f)(\sigma)\coloneqq f_{\sigma}$, i.e. the elements of $P$ get their functions values under $M(f)$ from the critical values of the corresponding critical cells.
Proposition 5.9.
The image of $\mathcal{MB}_{w}(X)$ is contained in $Mer_{wb}$.
Proof.
Since weak Morse–Benedetti functions cannot obtain the same value on different critical cells, the even stronger statement holds that $M(\mathcal{MB}_{w}(X))$ only contains injective maps for any regular CW complex $X$. ∎
We proceed by considering how crossing hyperplanes of $\mathcal{A}$ and $\mathcal{H}$ relates to edit moves between the induced combinatorial merge trees.
Lemma 5.10.
Let $X$ be a regular CW complex and let $H_{\sigma}^{\tau}$ be a hyperplane of $\mathcal{A}(X)$. If $\operatorname{dim}(\sigma)=0$, then passing from a Morse region $R_{+}$ on the positive side of $H_{\sigma}^{\tau}$ to a Morse region $R_{-}$ on the negative side without crossing any other hyperplane of $\mathcal{H}$ corresponds to the removal of a leaf. Moreover, passing through $H_{\sigma}^{\tau}$ the other way around corresponds to adding a leaf.
Remark 5.11.
We observe that a crossing as mentioned in Lemma 5.10 does not exist for all Morse regions on the positive side of $H_{\sigma}^{\tau}$. Even if it is possible to cancel $\sigma$ and $\tau$ without violating the Morse condition, it might happen that one has to cross other regions of $\mathcal{H}$ or even other Morse regions along the way.
Proof of Lemma 5.10.
Since, by assumption, both $R_{+}$ and $R_{-}$ are Morse regions and it is possible to pass from $R_{+}$ to $R_{-}$ without crossing any hyperplane of $\mathcal{H}$ other than $H_{\sigma}^{\tau}$, this means that for any Morse function $f\in R_{+}$, we have $f(\sigma)<f(\tau)$ and there is no cell with a function value between them. Since $\operatorname{dim}(\sigma)=0$, at filtration level $f(\sigma)$ a new connected component arises, which is merged with another connected component at level $f(\tau)$ via $\tau$. Because passing to $R_{-}$ corresponds to matching $\sigma$ and $\tau$, moving to $R_{-}$ removes the leaf of $M(f)$ that corresponds to $\sigma$. It is immediate that passing from $R_{-}$ to $R_{+}$ adds a new leaf to the induced merge tree. ∎
Remark 5.12.
It is well known, e.g. [For98, Theorem 11.1], that whenever a Morse function $f\in\mathcal{M}(X)$ induces a unique gradient path between a critical $d$ cell $\sigma$ and a critical $d+1$ cell $\tau$ such that all face relations along the gradient path are regular, the gradient path can be inverted, matching $\sigma$ and $\tau$ in the process. If $\sigma$ and $\tau$ have consecutive function values, then $\sigma$ is a face of $\tau$ and we are in the situation of Lemma 5.10. Otherwise, there are other cells with function values between $f(\sigma)$ and $f(\tau)$ and we have to cross hyperplanes of $\mathcal{A}$ corresponding to all face relations along said unique gradient path between $\sigma$ and $\tau$ as well as hyperplanes of $\mathcal{H}\setminus\mathcal{A}$ corresponding to cells that are not contained in the gradient path with function values between $f(\sigma)$ and $f(\tau)$. It should be remarked that during this procedure, we cannot cross all hyperplanes of $\mathcal{A}$ corresponding to the face relations along the unique gradient path simultaneously because otherwise we would leave the space of discrete Morse functions.
Lemma 5.13.
Let $X$ be a regular CW complex, let $\sigma,\tau\in X$ be two 1 cells, and let $H_{\sigma}^{\tau}$ be the corresponding hyperplane in $\mathcal{H}$. Let $f\in\mathcal{M}$ be a discrete Morse function inside a region $R_{1}$ of $\mathcal{H}$ such that $\sigma$ and $\tau$ are critical and merge connected components, i.e. correspond to inner nodes of $M(f)$. Assume there is a region $R_{2}$ adjacent to $R_{1}$ via $H_{\sigma}^{\tau}\in\mathcal{H}$, i.e. there is a path from $R_{1}$ to $R_{2}$ that only crosses $H_{\sigma}^{\tau}$ and no other hyperplane of $\mathcal{H}$. If
  1. (1)
    $\sigma$ and $\tau$ correspond to adjacent inner nodes of $M(f)$, then moving through $H_{\sigma}^{\tau}$ first merges the corresponding nodes and then splits them the other way around.
  2. (2)
    the nodes corresponding to $\sigma$ and $\tau$ are not comparable in $M(f)$, passing through $H_{\sigma}^{\tau}$ corresponds to moving their associated values past each other without changing the underlying combinatorial merge tree.
Remark 5.14.
In the context of Lemma 5.13, if the nodes that correspond to $\sigma$ and $\tau$ are comparable but not adjacent in $M(F)$, then it is not possible to cross $H_{\sigma}^{\tau}$ without crossing other hyperplanes of $\mathcal{H}$ because then there are other nodes with values in between due to the monotonicity of values associated to nodes of $M(f)$.
Proof of Lemma 5.13.
  1. (1)
    If $\sigma$ and $\tau$ correspond to adjacent inner nodes of the induced merge tree, then $\sigma$ and $\tau$ subsequently merge three previously disconnected connected components into one. Crossing $H_{\sigma}^{\tau}$ means that the components are merged simultaneously while inside $H_{\sigma}^{\tau}$ and the merging order is reversed after exiting $H_{\sigma}^{\tau}$ on the other side. This corresponds by definition to first merging the nodes and then splitting them the other way around.
  2. (2)
    If the nodes corresponding to $\sigma$ and $\tau$ are not comparable then $\sigma$ and $\tau$ belong to different connected components which are, if at all, merged at higher level via a third 1-cell. Hence, crossing $H_{\sigma}^{\tau}$ only changes the levels at which $\sigma$ and $\tau$ merge their respective components but not the underlying combinatorial merge tree.
Definition 5.15.
In order to topologize the space of merge trees $Mer$, we introduce the euclidean edit distance: We first define the metric for merge trees, whose underlying combinatorial merge trees are related by compositions of elementary edit morphisms (eem) in the same direction, i.e. either a composition of adding leaves and splitting inner nodes or removing leaves and merging inner nodes:
Let $\theta\in\operatorname{Map}_{\leq}(P,\mathbb{R}),\theta^{\prime}\in\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$ be merge trees such that $P$ and $P^{\prime}$ are related by a composition of eems in the same direction $\eta\colon P\rightarrow P^{\prime}$. Then by Remark 5.5 we have an induced morphism $\eta\colon P\rightarrow P^{\prime}$ if $\eta$ is a composition of adding leaves and splitting inner nodes. We extend the inclusion $\eta\colon P\rightarrow P^{\prime}$ associated to any composition of adding leaves and splitting inner nodes to a map $\eta_{*}\colon\operatorname{Map}_{\leq}(P,\mathbb{R})\times\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})\rightarrow\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$ by
(1) $\eta_{*}(\theta,\theta^{\prime})_{x}\coloneqq\begin{cases}\theta_{\eta^{-1}(x)}\text{ if }x\in\eta(P),\\ \theta_{\eta^{-1}(y)}\text{ if }x\notin\eta(P),\ x\text{ a leaf, }y\in\eta(P)\text{ is }x\text{'s parent in }P^{\prime}\\ \theta^{\prime}_{x}\text{ if }x\notin\eta(P)\text{ was added as a new parent node for a new leaf}\\ \theta^{\prime}_{y}\text{ if }x\notin\eta(P),\ x\text{ a leaf, }y\notin\eta(P)\text{ is }x\text{'s parent in }P^{\prime}\\ \theta^{\prime}_{y}\text{ if }x\notin\eta(P),\ x\text{ an inner node }y\text{ the node }x\text{ was split from}\end{cases}.$
If $\eta$ is a composition of removing leaves and merging inner nodes, we have an induced inclusion $\eta\colon P^{\prime}\rightarrow P$ in the other direction. Then we define
(2) $\operatorname{d}(\theta,\theta^{\prime})\coloneqq\operatorname{d}^{euc}(\eta_{*}(\theta,\theta^{\prime}),\theta^{\prime}).$
We define the euclidean edit distance by
(3) $\operatorname{d}(\theta,\theta^{\prime})\coloneqq\inf\limits_{\sigma}\{\sum\limits_{i}\operatorname{d}(\theta_{i},\theta_{i+1})\},$
where $\sigma=(\theta=\theta_{1},\dots,\theta_{n}=\theta^{\prime})$ are sequences of merge trees related by compositions of elementary edit morphisms in the same direction.
Proposition 5.16.
The euclidean edit distance $\operatorname{d}$ from Definition 5.15 is a pseudo metric on $Mer$ and a metric on $Mer_{<}$ and $Mer_{wb}$.
Proof.
The equality $\operatorname{d}(\theta,\theta)=0$ holds because the identity is an edit move. Symmetry holds because every edit move has an inverse: adding leaves is inverse to removing them and splitting inner nodes in inverse to merging them. Sequences of edit morphisms can be inverted stepwise by inverting elementary edit morphisms.
The triangle inequality follows from the triangle equality of the euclidean metric for trivial edit morphisms and by composition of the sequences in Equation 3 otherwise.
Let $\theta,\theta^{\prime}$ be merge trees over combinatorial merge trees $P,P^{\prime}$. If $P\cong P^{\prime}$, then positivity follows from the positivity of the euclidean metric. If $P\ncong P^{\prime}$ then $\operatorname{d}(\theta,\theta^{\prime})$ is by definition a sum of terms $\operatorname{d}(\theta_{i},\theta_{i+1})$ where $\theta_{i},\theta_{i+1}$ are merge trees over combinatorial merge trees $P_{i},P_{i+1}$ such that $P_{i}$ and $P_{i+1}$ are related by composition of elementary edit morphisms in the same direction. All these terms are non-negative because the euclidean metric is positive. Since $P\ncong P^{\prime}$, at least one of these terms must contain a non-trivial edit morphism. Let $\theta_{i},\theta_{i+1}$ be related by an non-trivial edit morphism. If $\theta_{i},\theta_{i+1}\in Mer_{<}$ then it follows from Equation 1 and Equation 2 that $\operatorname{d}(\theta_{i},\theta_{i+1})>0$ because values on new leaves/inner nodes coincide with the values of their parent nodes/nodes they are split from for $\eta_{*}(\theta_{i})$ but do not coinside for $\theta_{i+1}$. ∎
Theorem 5.17.
Let $X$ be a regular CW complex. Then the map $M\colon\mathcal{M}(X)\rightarrow Mer$ from Definition 5.8 that maps a discrete Morse function to its induced merge tree is continuous.
Proof.
Let $f_{0}\in\mathcal{M}(X)$ be an arbitrary discrete Morse function and let $\epsilon>0$. We have to show that there is a $\delta>0$ such that for all $f\in\mathcal{M}(X)$ with $\operatorname{d}(f,f_{0})<\delta$ we have $\operatorname{d}(M(f),M(f_{0}))<\epsilon$.
For that, we observe that for a fixed discrete Morse function $f_{0}$, the map to the induced merge tree is basically an orthogonal projection to the subspace of $\mathbb{R}^{X}$ which is spanned by the critical cells of dimensions 0 and 1 that correspond to the nodes of the induced merge tree. If $f_{0}$ lies in the interior of a region of $\mathcal{H}$, then we can choose $\delta$ small enough such that all $f$ with $\operatorname{d}(f,f_{0})<\delta$ lie in the same region of $\mathcal{H}(X)$ as $f_{0}$. Then all functions in the $\delta$ ball $U_{\delta}(f_{0})$ around $f_{0}$ induce the same underlying combinatorial merge tree and $M_{\lvert U_{\delta}(f_{0})}$ is just the orthogonal projection, which is continuous.
If $f_{0}$ lies on any of the hyperplanes of $\mathcal{H}$, then Lemma 5.10 and Lemma 5.13 describe how perturbations of $f_{0}$ affect the induced merge tree. That is, all merge trees in an $\epsilon$ neighborhood around $M(f_{0})$ that are induced by any discrete Morse function on $X$ are realized via a small perturbation of $f_{0}$. This is because merge trees in $U_{\epsilon}(f_{0})$ are related to $f_{0}$ by edit moves that correspond to crossing the hyperplanes that $f_{0}$ lies on. In this context, it is no problem that a priori $M(f_{0})$ admits more edit moves than $\mathcal{H}$ has hyperplanes because by Lemma 5.10 and Lemma 5.13, only the edit moves that correspond to a crossing a hyperplanes in the space of discrete Morse functions actually lead to merge trees that have preimages under the induced merge tree map. ∎
We define three hyperplane arrangements on the space of merge trees, which are related to the Morse arrangement and the braid arrangement on the space of Morse functions on any given CW complex:
Definition 5.18.
Let $P$ be a poset of merge tree type. We define the leaf arrangement $\mathcal{A}(P)$ in $\operatorname{Map}_{\leq}(P,\mathbb{R})$ by $\mathcal{A}(P)\coloneqq\{H_{x}^{y}|x\text{ is a leaf and }y\text{ is the parent node of }x\}$.
We define the order arrangement $\mathcal{O}(P)$ by $\mathcal{O}(P)\coloneqq\{H_{x}^{y}|x\leq y\}$.
Moreover, we define the braid arrangement $\mathcal{H}(P)$ in $\operatorname{Map}_{\leq}(P,\mathbb{R})$ by $\mathcal{H}(P)\coloneqq\{H_{x}^{y}|x,y\in P\}$.
Proposition 5.19.
Let $P$ be a poset of merge tree type and let $f\colon X\rightarrow\mathbb{R}$ a discrete Morse function on a CW complex $X$ such that $P$ is isomorphic to the underlying combinatorial merge tree of $M(f)$. Then the following holds:
  1. (1)
    The hyperplanes of $\mathcal{A}(P)$ correspond to families of maximal gradient paths between critical 0 and 1 cells induced by $f$. If $f$ is weakly Morse–Benedetti, then this correspondence is bijective.
  2. (2)
    The hyperplanes of $\mathcal{O}(P)$ correspond to a collection of zig-zags of maximal gradient paths between cells of dimension $\leq 1$ induced by $f$ that begin and end at maximal critical cells of sublevel complexes without leaving the sublevel complex corresponding to the higher of the two cells.
  3. (3)
    The hyperplanes of $\mathcal{H}(P)$ correspond to a collection of zig-zags of maximal gradient paths between cells of dimension $\leq 1$ induced by $f$.
In particular $\mathcal{A}(P)\subset\mathcal{O}(P)\subset\mathcal{H}(P)$ holds.
Proof.
  1. (1)
    By Definition 5.8, the leaves of $P$ correspond to critical 0 cells of $f$. Moreover, the parent relation $x\prec y$ of any leaf $x$ corresponds to the collection of critical 1 cells that merge the connected component that corresponds to $x$, say $X_{M(f)_{y}-\varepsilon}[\sigma]$, where $\sigma$ is the unique critical 0 cell, with the connected components that correspond to $y$’s other child nodes. Let $\tau$ be one of these merging 1 cells. Then one of $\tau$’s boundary 0 cells $\sigma_{0}$ belongs to $X_{M(f)_{y}-\varepsilon}[x]$ and there is a gradient path $\gamma$ in $X_{M(f)_{y}}[\tau]$ that begins with $\tau\supset\sigma_{0}$. Since maximal gradient paths lead to critical cells and cannot lead through cells of increasing dimension, $\gamma$ has to lead to $\sigma$. If $f$ is weakly Morse–Benedetti, then each inner node $y$ corresponds to a unique merging 1 cell $\tau$ and because, by assumption, only one of $\tau$’s boundary 0 cells $\sigma_{0}$ belongs to $X_{M(f)_{y}-\varepsilon}[x]$, the gradient path starting at $\tau\sup\sigma_{0}$ is unique because it is contained in the 1 skeleton of $X$.
  2. (2)
    The hyperplanes of $\mathcal{O}(M(f)$ correspond by definition to inclusions of connected components sublevel complexes of $f$. Similar to the proof of (1), cover relations $x\prec y$ of nodes in $M(f)$ correspond to gradient paths from merging 1 cells that correspond to $y$ to any critical 0 cell of the connected component that corresponds to $x$. If $x\leq y$ is not a cover relation, then the statement follows inductively because the interval $[x,y]$ is a chain due to $P$ being of merge tree type.
  3. (3)
    Since posets of merge tree type have a unique maximum, namely the root, by definition, all elements of $P$ are related by zig-zags of comparison relations. It follows from (2) that cells that correspond to elements of $P$ are related by zig-zags of zig-zags of gradient paths, which are zig-zags of gradient paths.
Remark 5.20.
Since discrete gradient paths are sequences of matched cells, gradient paths correspond to intersections of hyperplanes of $\mathcal{A}(X)$. Even though this observation makes the correspondence between families of gradient paths and hyperplanes in Proposition 5.19 (1) into a correspondence between certain intersections of hyperplanes in $\mathcal{A}(X)$ and hyperplanes in $\mathcal{A}(M(f)$, that does not in general lead to a correspondence between cancellations of gradient paths and removals of leaves.
Corollary 5.21.
If $f\colon X\rightarrow\mathbb{R}$ is a weak Morse–Benedetti function, then cancellations of critical 0 cells and 1 cells are in 1-1 correspondence with removals of leaves of the induced merge tree. If $f$ is not weakly Morse–Benedetti, cancellations of critical 0 cells and 1 cells still induce removals of leaves of the induced merge tree but not necessarily the other way around.
Corollary 5.22.
Let $P$ be a poset of merge tree type such that $P\cong M(f)$ for a weak Morse–Benedetti function $f\colon X\rightarrow\mathbb{R}$, then $M^{-1}(\mathcal{A}(P))\subset\mathcal{A}(X)$. Moreover, $M^{-1}(\mathcal{A}(P))$ agrees with $\mathcal{A}(X)$ on the 0 skeleton of $X$.
5.3. The space of barcodes
We identify sets of barcodes with sets of order-preserving maps $\operatorname{Map}_{\leq}(P,\mathbb{R})$ for posets $P$ of barcode type.
Definition 5.23.
We define the set of barcodes by
$Bar\coloneqq\coprod\limits_{P\text{ of barcode type}}\operatorname{Map}_{\leq}(P,\mathbb{R})$
Remark 5.24.
Technically, we have to solve the same set theoretic issues as for the space of merge trees again for the space of barcodes. They can be solved in the same way as for merge trees.
Definition 5.25.
Let $P,P^{\prime}$ be two posets of barcode type.
  1. (1)
    If $P\cong P^{\prime}$ as posets, we call an isomorphism $\eta\colon P\rightarrow P^{\prime}$ a reordering of the bars. We call $\operatorname{id\colon P\rightarrow P}$ the trivial reordering.
  2. (2)
    We say that $P^{\prime}$ arises from $P^{\prime}$ by collapsing a bar if $P^{\prime}$ has one bar fewer than $P$. In that case, a collapse of a bar is an injective poset map $\eta\colon P^{\prime}\rightarrow P$. Moreover, we say that $P$ arises from $P^{\prime}$ by creation of a new bar.
We refer to all of the mentioned operations as edit moves. An elementary edit morphism between posets of barcode type is a composition of an edit move with a reordering of the bars. An edit morphism is a composition of elementary edit morphisms.
Associated to an edit morphism $\eta\colon P\rightarrow P^{\prime}$, we define a map $\eta_{*}\colon\operatorname{Map}_{\leq}(P,\mathbb{R})\times\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})\rightarrow\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$ by
$\eta_{*}(\beta,\beta^{\prime})_{x}\coloneqq\begin{cases}\theta_{x}\text{ if }x\in P,\\ \theta^{\prime}_{x}\text{ if }x\notin P\end{cases}.$
Definition 5.26.
The euclidean edit distance on $Bar$ is defined as follows:
Let $P,P^{\prime}$ be posets of barcode type. Let $\beta_{1}\in\operatorname{Map}_{\leq}(P,\mathbb{R}),\beta_{2}\in\operatorname{Map}_{\leq}(P^{\prime},\mathbb{R})$. If $P^{\prime}$ arises from $P$ by an edit morphism $\eta\colon P\rightarrow P^{\prime}$, we define $\operatorname{d}(\beta_{1},\beta_{2})\coloneqq\operatorname{d^{euc}(\eta_{*}(\beta,\beta^{\prime}),\beta^{\prime})}$ In the general case, we define the euclidean edit distance by
$\operatorname{d}(\beta,\beta^{\prime})\coloneqq\inf\limits_{\sigma}\{\sum\limits_{i}\operatorname{d}(\beta_{i},\beta_{i+1})\},$
where $\sigma=(\beta=\beta_{1},\dots,\beta_{n}=\beta^{\prime})$ are sequences of barcodes related by edit morphisms.
We topologize the space of barcodes with the euclidean edit distance.
Definition 5.27.
Let $P$ be a poset of barcode type. We define the order arrangement $\mathcal{\mathcal{O}(P)}$ in $\operatorname{Map}_{\leq}(P,\mathbb{R})$ by $\mathcal{O}(P)\coloneqq\{H_{x}^{y}|x\leq y\}$.
Moreover, we define the braid arrangement $\mathcal{H}(P)$ in $\operatorname{Map}_{\leq}(P,\mathbb{R})$ by $\mathcal{H}(P)\coloneqq\{H_{x}^{y}|x,y\in P\}$.
In order to define a map from the space of strict merge trees to the space of barcodes, we apply the combinatorial elder rule, see [${}^{+}$24, Definition 2.15]
Definition 5.28.
We define a map $B\colon Mer_{wb}\rightarrow Bar$ as follows:
Let $\theta\in Mer_{<}$ be a strict merge tree with an underlying combinatorial merge tree $P$. We define a poset $\tilde{B}(\theta)$ inductively: Add for each inner node $y$ of $P$ an interval $[x,y]$ to $\tilde{B}(\theta)$, where $x$ denotes leaf of $P_{\leq y}$ with the minimal value $\theta_{x}$. For all leaf nodes $x$ that are not matched by the rule above, we add an interval $[x,y]$ to $\tilde{B}(\theta)$ where $y$ is $x$’s parent node.
By construction, we have a canonical map $i\colon\tilde{B}(\theta)\rightarrow P$ given by mapping the elements of $\tilde{B}(\theta)$ to their corresponding nodes in $P$. We define $B(\theta)\coloneqq i^{*}(\theta)=\theta\circ i$.
Proposition 5.29.
The map $B\colon Mer_{wb}\rightarrow Bar$ from Definition 5.28 is continuous.
Proof.
Let $\theta_{0}\in Mer_{wb}$ be an arbitrary well-branched merge tree and let $\epsilon>0$. We have to show that there is a $\delta>0$ such that for all $\theta\in Mer_{wb}$ with $\operatorname{d}(\theta,\theta_{0})<\delta$ we have $\operatorname{d}(B(\theta),B(\theta_{0}))<\epsilon$.
Since $\theta_{0}$ is by assumption strict, it must lie in the interior of some region of $\mathcal{O}(s(\theta))$. Thus, $\delta$ can be chosen small enough such that all $\theta$ with $\operatorname{d}(\theta,\theta_{0})<\delta$ have the same underlying combinatorial merge tree $s(\theta)$ and induce the same combinatorial barcode $\tilde{B}(\theta)$. Therefore, $B_{\lvert U_{\delta}(\theta_{0})}$ is a linear map between finite-dimensional vectorspaces and, therefore, continuous. ∎